UT 20160530 LUCI AO Commissioning Interrupt

Executing 01_WunitTransform

Handover from observers 12:48 MST (07:48 UT)

Config both sides with N30_FieldStop mask, BrG, Mirror, FC off.

Darks:

0001 0003 LUCI1 3x2.75s LIR:INT

0133 0135 LUCI2 3x2.75s LIR:INT

Instrument configs threw errors related to the CU lamps

...init both units (did that earlier, though)

...aaaaand now we have counds! Where did they come from?

08:50 Closing. Seeing blew up and clouds still here.

LUCI1 Field Stop adjusted

LUCI2 Field Stop needs to be moved left 280 pix, up 270 pix before next interrupt

Error committing instrument setups...INCORRECT_CALIBRATION_LAMP_STATE format

The LAMP block is missing from the script? Corrected

- Does this lamp block need to be in each obsItem with a preset in it?

### OSA Night Log:

Created with: LBTO OSA Night Report Tool v6.3
Submitted by: John Hill < [[mailto:jhill@lbto.org][jhill@lbto.org]]>
Submitted on: Mon May 30 10:54:02 2016

===========================================================================

UTdate:  2016-05-30


ON DUTY:
------------------------
OSA:  John Hill
ISA:  Barry Rothberg
ISp:  John Morris
Tel:  Michael Wagner
Soft: Michele De La Pena
AO:   Greg Taylor
Eng:  Tom McMahon


OBSERVERS:
------------------------
Partner: LBTB - Germany
Names:   Alessandra Contursi
 Sirio Belli


INSTRUMENTS:
------------------------
L_Start: 
R_Start: LUCIFER
L_End:   
R_End:   


CONDITIONS:
------------------------
Evening:  O:Photometric
Eve_IQ:   0.8
Midnight: O:Spectroscopic
Mid_IQ:   0.7
Morning:  X:Clouds
Morn_IQ:  


TIMES:
------------------------
UThandover: 00:27
UTopened:   02:20
UTclosed:   08:55
UTSciStart: 03:23
UTSciEnd:   11:11


ALMANAC:
------------------------
For the night of: Sun, 2016 May 29 ---> Mon, 2016 May 30
 Local midnight = 2016 May 30,  7 hr UT, or JD 2457538.792
 Local Mean Sidereal Time at midnight =  16 13 20.5
 Sunset ( 1000 m horizon):   19 27 MST; Sunrise:   5 08 MST
 Evening twilight:  21 00 MST;  LMST at evening twilight:  13 13
 Morning twilight:   3 34 MST;  LMST at morning twilight:  19 48
 12-degr twilight: 20 23 MST -->  4 11 MST; night center:   0 17 MST
 Moonrise:   1 17 MST   
 Moon at civil midnight: illuminated fraction 0.408
 0.8 days after last quarter, RA and dec:  23 24 14,  -4 49.3
 The sun is down for  9.7 hr;  6.6 hr from eve->morn 18 deg twilight.
  4.3 dark hours after end of twilight and before moonrise.


DOWNTIME:
------------------------
03:50to03:58 = 00:08 > Human_Cause.PI > Guide star for blind offset moves out of field
07:48to08:06 = 00:18 > Reconfiguration > Reconfigure and authorize for LUCI/LUCI
08:10to08:55 = 00:45 > Weather > Heavy clouds prevent useful work.
08:55to11:11 = 02:16 > Weather > Closed for heavy clouds.
------------------------
07:48 Length of Night (12deg-12deg Twilight)
03:27 Total Downtime  (Series)
03:27 Total Downtime  (Parallel)
55.8% On-Sky Efficiency


===========================================================================


SUMMARY
------------------------

Beautiful clear sky with just a touch of smoky haze, and good seeing.

Utility Building Chiller problems in the early evening.

Switched to LUCI/AO Interupt at 07:48 which got nowhere as clouds moved in.

Closed for heavy clouds at 08:55, never reopened.


LOG
------------------------

23:50UT rest shells and move to horizon for LBC fills
DIMM apertures are 70 mm diameter, and are spaced 300 mm apart (center-center).

00:25UT to zenith, 120 for LUCI calibrations (HBS, AZ, EL, ENC still running)
00:33UT LUCI2 MOS makes a temperature alarm (apparently a sensor glitch)
00:36UT ersetpoint 0; elloadfromabs, azloadfromabs
01:03UT homed agw2; agw1, SX M3, SX M2, DX M2, DX M3
01:19UT Tam=10.5 Tch=9.8 Tgl(ref):SX=6.2(9.3),DX=5.9(9.4) Hum=18.4% Wind=6.8@185\xB0
01:20UT ZeroRE on both primaries; az/elencsettype strip

01:38UT SX AdSec IDL unhappy about still getting slopes. adsc_stop, adsc_start (slope counter was not incrementing)
02:00UT Start mirror ventilation both sides
02:04UT Authorized: NONE / LUCIFER
02:15UT Test ACQUIRE preset RA 11:49:03.9 DEC +14:34:20.4 AZ 100.0 EL 31.0 Acquisition

02:17UT Shutter doors open
02:26UT Rear vent doors open 25%.
02:26UT Tam=9.2 Tch=9.8 Tgl(ref):SX=6.6(9.2),DX=6.4(9.2) Hum=20.0% Wind=7.5@229\xB0
02:28UT ACQUIRE preset RA 11:49:03.9 DEC +14:34:20.4 AZ 173.1 EL 71.7 Acquisition
02:32UT Pointing model change - IE/CA: DX => 24.0/-238.0
02:36UT ACQUIRE preset RA 10:26:30.2 DEC +59:09:12.0 AZ 338.5 EL 60.3 Acquisition
02:43UT Pointing model change - IE/CA: DX => 14.0/-253.0
02:57UT John was trying different binnings on the guider while waiting for twilight.
It worked OK at bin4 and bin1, but graphic overlays on GUI lag the guide image by
a frame.
03:00UT ACTIVE preset RA 10:43:43.3 DEC +57:11:57.1 AZ 334.5 EL 61.2 Standard
03:01UT Telluric Standard
03:02UT Tam=8.7 Tch=9.3 Tgl(ref):SX=7.1(8.8),DX=6.9(8.8) Hum=20.0% Wind=6.0@229\xB0
03:02UT SEEING: DIMM 0.99" @63.4\xB0, GCSR 0.81" (LBT @61.1\xB0)
03:08UT reduce rear vent doors to 15%
03:35UT Utility Chiller Problem - return is at +9 degC send Mike to investigate
03:35UT Stop Both Mirror Ventilation to shed load
03:39UT ACTIVE preset RA 14:19:39.5 DEC +52:52:33.6 AZ 391.4 EL 64.5 Science
Science target
03:42UT Pointing model change - IE/CA: DX => 12.0/-255.0
03:42UT Tam=8.4 Tch=9.1 Tgl(ref):SX=7.6(8.6),DX=7.5(8.6) Hum=20.1% Wind=5.4@234\xB0
03:43UT SEEING: DIMM 0.98" @71.0\xB0, GCSR 0.67" (LBT @64.9\xB0)
03:50UT Blind offset moved guide star out of the available range of the probe. Is it downtime if we are taking unguided spectra?
03:58UT ACTIVE preset RA 14:19:39.5 DEC +52:52:33.6 AZ 386.3 EL 66.4 Science
w/ Different Guide Star
03:59UT Mike has revived the chiller, so restart both Mirror Ventilation
04:42UT Some thin clouds have appeared in the SW ruining the photometricity (no, it isn't a word). It is still clear at our target in the north.
04:43UT Tam=8.0 Tch=8.5 Tgl(ref):SX=7.8(8.3),DX=7.7(8.1) Hum=20.7% Wind=8.0@237\xB0
04:43UT SEEING: DIMM 0.93" @73.7\xB0, GCSR 0.66" (LBT @69.5\xB0)
04:52UT fully close REAR VENT DOORS as wind is getting up to 10 m/sec.
05:40UT reopen REAR VENT DOORS to 15%.
05:40UT Tam=7.2 Tch=8.0 Tgl(ref):SX=7.9(8.0),DX=7.7(7.6) Hum=21.2% Wind=8.5@242\xB0
05:40UT SEEING: DIMM 0.99" @69.7\xB0, GCSR 0.69" (LBT @69.0\xB0)
06:48UT Tam=6.4 Tch=7.2 Tgl(ref):SX=7.7(7.5),DX=7.4(7.2) Hum=23.6% Wind=7.8@215\xB0
06:48UT SEEING: DIMM 1.10" @60.5\xB0, GCSR 0.60" (LBT @62.9\xB0)c!

07:48UT INTERUPT START FOR LUCI/AO
07:48UT Move to zenith; rest shells; insert stow pins
07:51UT Deploy SX M3 swing arm
07:52UT Authorize LUCI/LUCI
08:07UT ACQUIRE/ACQUIRE preset RA 17:34:56.0 DEC +12:33:38.1 AZ 169.6 EL 69.6 Acquisition
08:07UT Authorized: LUCIFER / LUCIFER
08:07UT Pointing model change - IE/CA: SX => -11.0/-32.0; DX => 29.0/-248.0
08:08UT ACE_ACTIVE/ACE_ACTIVE preset RA 18:03:47.7 DEC -00:17:23.7 AZ 160.8 EL 55.5 Collimation
The LUCI AO guys (Dave and Doug) must have bad karma - suddenly we have 4+ mags of clouds moving in from the south.
08:10UT Lost to weather start
08:15UT Pointing model change - IE/CA: SX => -11.0/-29.0; DX => 29.0/-248.0
08:25UT Pointing model change - IE/CA: SX => -15.0/-22.0; DX => 27.0/-248.0
08:26UT ACTIVE/ACTIVE preset RA 17:53:03.3 DEC -02:34:45.6 AZ 174.1 EL 54.6 Collimation gs=14
08:28UT ACE_ACTIVE/ACE_ACTIVE preset RA 18:03:47.7 DEC -00:17:23.7 AZ 169.8 EL 56.6 Collimation
08:37UT Close REAR VENT DOORS
08:39UT ACE_ACTIVE/ACE_ACTIVE preset RA 18:03:47.7 DEC -00:17:23.7 AZ 174.8 EL 56.9 Collimation
08:47UT ACTIVE/ACTIVE preset RA 17:44:15.8 DEC -00:02:16.1 AZ 187.4 EL 57.0 Collimation Stone_M1
08:48UT Clear Active Optics on SX M1
08:54UT Slew to 31 deg to close
08:55UT Closing shutter doors
08:59UT Move to zenith - SX AdSec is in a panic state. DougM fixed it.
09:10UT Authorized: NONE / LUCIFER
09:09UT Allsky camera waited almost 5 minutes between images. Why?
10:43UT Park telescope at 120.6,90, waitopr, enclosure off, HBS off, exit DIMM, park rotators


-- %USERSIG{DavidThompson - 2016-06-02}%

Comments


 
Topic revision: r2 - 06 Jun 2016, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback